Let f(x, y) = y cos (5x²) + 3x²644 Find fxy? (calculus 3)
Answer #1

Given the function
f(x,y) = y cos(5.) + 3.r-ey
And so, taking a partial derivative w.r.t. x, we get
fr(2,y) = af (2,y) ar a =a (y cos(5.rº) + 3c2e4y)
  f (x,y) = -10.ry sin(5x) + 6.cey
And now, taking partial derivative w.r.t. y, we get

  Say(x, y) = O520, y) = 2 (-10-y sin(5r?) + 6ee4u )
  fry(x,y) = -106 sin(5.c) + 24.cey
Another way, you can first take partial derivative w.r.t. y and then take partial derivative w.r.t. x. But, you will get the same answer either way, because of the fact that fry(,y) = fyr (1,7) .